# Phase Transitions in Anisotropic Turbulence

### v1

@@ III Dimensional transitions
In view of the strikingly disparate properties of turbulence in two and three spatial dimensions, it is natural to examine what has been termed the dimensional transitionDeusebio et al. (2014); Boffetta (2023) between these two cases as the degree of anisotropy is increased starting from an isotropic 3D turbulent flow, e.g., by flattening the domain geometry. It is interesting to note in this context that in three dimensions, since the only quadratic invariant in addition to energy is helicity (cf. Sec. II.1), which is sign indefinite and not related to the energy, there is no Fjørthoft-type argument analogous to the 2D case. This indicates that although in homogeneous and isotropic 3D turbulence energy does cascade to small scales, this is not necessarily true in strongly anisotropic 3D flows. Indeed, the dimensional transition typically occurs in two steps: first, as the flow becomes quasi-2D, the forward energy cascade splits up, with some energy cascading to large scales, while 3D variations persist in the flow, transferring the remainder of the energy to small scales. Provided $L\ll L_{\alpha}$, even a small inverse energy flux can lead to condensation at late times, thus drastically altering the flow. Second, for very strong anisotropy, 3D variations decay and exact 2D flow is obtained (forcing and boundary conditions permitting).
Below, we briefly summarize important known results on both on these steps of the dimensional transition from 3D to 2D turbulence in different systems.
@@ III.1 Thin-layer turbulence
where $\overline{(\cdot)}$ represents the vertical average.
@@ III.1.1 Homogeneous quasi-2D turbulence
The first work showing that energy can cascade simultaneously to large and small scales in turbulence confined within thin layers was made by Smith et al.,Smith, Chasnov, and Waleffe (1996) and a number of more detailed subsequent works later quantified energy fluxes as a function of the layer height.Celani, Musacchio, and Vincenzi (2010); Musacchio and Boffetta (2017); Benavides and Alexakis (2017); van Kan and Alexakis (2019); Poujol, van Kan, and Alexakis (2020) Below, based on these works, we describe how $D_{\nu}^{\mathcal{E}},D_{\alpha}^{\mathcal{E}}$ $\mathcal{E}_{LS}$ and $\mathcal{E}_{3D}$ change as the layer height $H$ is varied.
@@
where an exponent $\beta_{1}$ close to unity is observed, but more work is required to measure it more precisely and complement the existing numerical evidence.Benavides and Alexakis (2017); van Kan and Alexakis (2019) No theory exists to describe the physics of this non-equilibrium transition, and therefore more detailed numerical and theoretical studies will be needed to clarify its properties and determine whether it may potentially be related to a known universality class. The very existence of a critical point at $H=H_{3D}$ is nontrivial, and remains to be further investigated, given that even bifurcations in extremely well-controlled low-dimensional dynamical systems show some rounding due to experimental imperfections. While the existence of $H_{3D}$ has been shown to be robust,Poujol, van Kan, and Alexakis (2020) whether the forcing acts on 2D modes alone or a mixture of 2D and 3D modes, more work is needed to further clarify the impact of the forcing dimensionality on this transition.
@@
where $\eta=\epsilon k_{f}^{2}$ is the enstrophy flux, although this remains to be tested in full direct numerical simulations. The energy reaching the scale $H$ is then transported, by 3D interactions, to even smaller scales.
@@
Figure 3: Visualization of $\omega_{z}^{2}$ in typical flow states of thin-layer turbulence, showing a top view and a side view of the layer. Left panel: Flow in a deep layer at $H>H_{\rm 3D}$ (see main text). Right panel: large-scale vortex condensate in a thin layer at $H<H_{\rm 3D}$. Figure adapted from van Kan et al. (2019)van Kan and Alexakis (2019).
@@ III.1.2 The quasi-2D condensate regime
Next, we consider the case where the bottom drag is small or absent. In a finite domain, even a small inverse energy flux then leads to the formation of a large-scale condensate. While numerical simulations of large-scale condensates are computationally very demanding since the saturation of the condensate amplitude is a slow process requiring very long integration times, a few numerical studies exist in the literaturevan Kan and Alexakis (2019); Musacchio and Boffetta (2019) which explicitly address condensate formation in thin-layer turbulence. Laboratory experiments, on the other hand, do not suffer from the same restrictions in terms of observation time as simulations, although they are more limited when it comes to varying the layer geometry or the forcing scale, and a number of experimental studies of condensates in thin-layer turbulence have been reported,Xia, Shats, and Falkovich (2009); Shats, Byrne, and Xia (2010); Xia et al. (2011); Byrne, Xia, and Shats (2011); Xia and Francois (2017) although these have not yet been able to explore the layer height dependence in the same detail.
As in the above discussion of the homogeneous turbulence case, we begin by considering a deep layer of height $H\gg\ell_{f}$, where a forward energy cascade is present and no condensate forms. As the layer height is reduced to the critical height $H=H_{3D}$ (discussed above), a weak inverse cascade emerges leading to the formation of a condensate with large-scale energy $\mathcal{E}_{LS}$. It was shownvan Kan and Alexakis (2019) (for the case $\alpha=0$) that the transition to the condensate is discontinuous. Specifically, for $H$ slightly larger than $H_{3D}$, no condensate formed when initializing the flow at rest, and the large-scale $\mathcal{E}_{LS}$ was negligible. By contrast, for $H$ just below $H_{3D}$, the large-scale energy $\mathcal{E}_{LS}$ jumped to a finite value, taking up nearly all of the kinetic energy of the flow, indicating the presence of the condensate. Furthermore, the transition was shown to be hysteretic: as $H$ was increased again, the condensate was found to persist at values of $H$ where it did not spontaneously emerge from small-amplitude initial conditions within the integration time. In other words, there is a coexistence of two attractors (illustrated in Fig. 3 in terms of $\omega_{z}^{2}$) in a range of $H$ near $H_{3D}$. When simulating the system within this range for very long times, one observes rare spontaneous transitions between the two attactors, induced by turbulent fluctuations, occurring at random times. This phenomenon will be described in more detail in section IV.
@@
As in the homogeneous quasi-2D turbulence case, when the layer height is reduced to the order of the Kolmogorov scale, another phase of turbulence is encountered where 3D perturbations decay viscously. However, the dynamics of 3D modes at $H$ slightly above $H_{2D}$ are strongly impacted by the presence of the condensate. In contrast with homogeneous quasi-2D turbulence, where 3D perturbations can develop anywhere in the flow, the condensate flow is rendered highly inhomogeneous by the presence of large-scale coherent structures, with larger strain and vorticity in some regions than others, while the flow still remains chaotic. The evolution of infinitesimal 3D perturbations on the background of a turbulent 2D condensate was recently investigated by Seshasayanan et al. in direct numerical simulations,Seshasayanan and Gallet (2020); Lohani, Nayak, and Seshasayanan (2024) revealing that $\mathcal{E}_{3D}$ in this limit exhibits a random behavior with long periods of slow viscous decay and short periods of rapid exponential increase. The periods of increase were found to appear when extremes of the vorticity or the strain rate crossed a certain threshold. Considering 2D point vortex flow coupled to localized 3D perturbations, van Kan et al.van Kan, Alexakis, and Brachet (2021a) subsequently showed that the instantaneous growth rate of the energy of the 3D perturbations displayed power law distributions that were linked to the power law distribution of strain in space. This indicates that the logarithm of the energy of the 3D perturbations undergoes Lévy random motion, explaining the sudden jumps in the growth of the perturbation energy observed in the simulations of Seshasayanan et al. This led to the investigation of this type of random process in simplified stochastic dynamical systems, where a new type of intermittency was identified (distinct from the dissipative intermittency of 3D turbulence mentioned in Sec. II), so-called Lévy on-off intermittency.van Kan, Alexakis, and Brachet (2021b); Van Kan and Pétrélis (2023) In particular, this theory predicts a distinct power-law behavior of the 3D energy
@@ III.2 Rapidly rotating turbulence
which was also verified in the limit $\rm Ro\to 0$ using a reduced, asymptotically valid set of equations.van Kan and Alexakis (2020)
@@
Stably stratified turbulence, or stratified turbulence for short, refers to turbulence under the influence of gravity $-g\hat{\mathbf{z}}$ and an imposed, stable background density gradient $S=-\rho_{0}{\rm d}\rho(z)/{\rm d}z>0$, where $\rho_{0}$ is a reference density. Like rotating flow, stratified flow supports wave motion, namely internal waves whose frequency depends on stratification. Unlike rotation, however, stratification acts to suppress the formation of 2D turbulence: by inhibiting vertical motion due to the energy cost associated with lifting up heavy fluid parcels, it leads to the formation of thin layers, creating strong vertical gradients in the process. The strength of stratification is measured by the nondimensional Froude number ${\rm Fr}=\epsilon^{1/3}/(gS)^{1/2}/\ell_{f}^{2/3}$, where small values of $\rm Fr$ indicate strong stratification. Sozza et al.Sozza et al. (2015) showed that the critical height, at which the bidirectional energy cascade emerges, decreases with increasing stratification strength as
@@ III.5 Convection
Convection refers to fluid motion induced by buoyancy in the presence of an imposed unstable background density gradient ($S<0$). The paradigmatic model of this process, with the fluid being contained between a hot bottom boundary and a cold top boundary, is Rayleigh-Bénard convection. Convection induces three-dimensional vertical motion. The resulting dynamics are highly nontrivial, in particular since the buoyant forcing is self-organized in response to the flow, rather than being an externally imposed body force. Despite these differences, it has been found that convection, under highly anisotropic conditions, can also display quasi-2D behavior and the formation of large-scale flow structures.
The most well-studied example is that of rapidly rotating convection.Julien et al. (2012a); Ecke and Shishkina (2023) There, for sufficiently small Rossby numbers, an inverse energy cascade forms and results in the formation of large-scale vortices or jets,Guervilly, Hughes, and Jones (2014); Favier, Silvers, and Proctor (2014); Rubio et al. (2014); Guervilly and Hughes (2017); Julien, Knobloch, and Plumley (2018); Maffei et al. (2021) not unlike the systems discussed in the previous sections. Furthermore, the behavior of the large-scale vortex condensates also resembles thin-layer and rotating turbulence, since they emerge via a discontinuous transition.Favier, Guervilly, and Knobloch (2019); de Wit et al. (2022) However, the case of convection has additional complexity, in particular due to the presence of thermal and momentum boundary layers. Recent progress in numerically simulating convection at very low Rossby numbers has furthermore revealed a phase transition between fully rotationally constrained geostrophic turbulence and a regime where boundary layers exhibit ageostrophic turbulence,van Kan et al. ; Julien et al. which had been predicted theoretically.Julien et al. (2012b) It is known from direct numerical simulations of rotating convection that, in certain parameter regimes, large-scale vortices form only for stress-free boundary conditions on top and bottom, but are suppressed with no-slip boundary conditions on the top and bottom walls.Kunnen et al. (2016) Importantly, the corresponding problem of the robustness of inverse cascades and condensation with respect to no-slip boundaries, which are typically present in laboratory experiments, remains to be studied in mechanically forced thin-layer turbulence.
@@ IV.1 Reversals of mean fields
Spontaneously reversing large-scale fields are commonly found in turbulent fluid flows. An important example is the Earth’s magnetic field, which is thought to arise from the dynamo effect Fauve and Petrelis (2003) associated with motions in the liquid metal outer core, and whose polarity is known to reverse at random times.Jacobs (1994) By contrast, the Sun’s magnetic field follows the more regular 22-year solar cycle.Charbonneau (2014) Magnetohydrodynamic dynamo reversals have also been reproduced in laboratory experiments,Berhanu et al. (2007) and were analyzed extensively at the theoretical level.Pétrélis et al. (2009); Gallet et al. (2012) An additional example of a reversing flow is found at high altitudes in the Earth’s atmosphere, where the so-called quasi-biennial oscillation consists in reversals of the mean zonal wind approximately every two years,Baldwin et al. (2001) a phenomenon that has also been studied using laboratory analogs.Plumb and McEwan (1978); Semin et al. (2018) In addition, reversals are also common in convectively driven flows and have been studied in detail in that context.Sugiyama et al. (2010); Ni, Huang, and Xia (2015); Liu et al. (2024)
Another setting where random reversals are observed is that of confined, large-scale, 2D or quasi-2D turbulent flows, which have been studied both in the laboratorySommeria (1986); Michel et al. (2016); Fauve et al. (2017); Pereira, Gissinger, and Fauve (2019) and in theoretical work.Mishra et al. (2015); Shukla, Fauve, and Brachet (2016); Dallas, Seshasayanan, and Fauve (2020a) Shukla et al.Shukla, Fauve, and Brachet (2016) introduced the approach of investigating the statistics of the large-scale modes based on the truncated Euler equation of ideal fluids, retaining only a finite number $N$ of Fourier modes with wave numbers below a chosen cutoff. Such truncations naturally arise in any standard pseudospectral PDE solver. They considered the flow inside a square domain $[0,\pi]^{2}$ with free-slip boundary conditions, where the stream function can be expanded as $\psi(x,y)=\sum_{n,m}\hat{\psi}_{n,m}\sin(nx)\sin(my)$, with the sum being over a truncated set containing $N$ modes. Shukla et al. compared direct numerical simulations of the 2D Navier-Stokes equation for a viscous fluid with corresponding solutions of the 2D truncated Euler equation and were able show good agreement between the statistics in the two cases. Similar findings were also made in a forced 2D shear flow.Dallas, Seshasayanan, and Fauve (2020b)
@@
Figure 5: Left panel: Spontaneous reversals seen in the time series of the large-scale mode $\hat{\psi}_{1,1}$ in a confined 2D flow described by the Euler equation truncated to $N=13$ modes, obtained from numerical simulations. Right panel: probability density function (PDF) $p(\hat{\psi}_{1,1})$ versus large-scale circulation $\hat{\psi}_{1,1}$ with different colored lines representing different values of the control parameter $\delta=\Omega/(k_{2}^{2}\mathcal{E})-1$ ($\delta>0$ increases from bottom to top). The black dashed lines indicate the theoretically predicted form of $p(\hat{\psi}_{1,1})$ at different values of $\delta$, which agrees well with the numerical result. Figure adapted from van Kan et al. (2022).van Kan, Alexakis, and Brachet (2022)

### v2

@@ III Dimensional transitions
In view of the strikingly disparate properties of turbulence in two and three spatial dimensions, it is natural to examine what has been termed the dimensional transitionDeusebio et al. (2014); Boffetta (2023) between these two cases as the degree of anisotropy is increased starting from an isotropic 3D turbulent flow, e.g., by flattening the domain geometry. It is interesting to note in this context that in three dimensions, since the only quadratic invariant in addition to energy is helicity (cf. Sec. II.1), which is sign indefinite and not related to the energy, there is no Fjørtoft-type argument analogous to the 2D case. This indicates that, although in homogeneous and isotropic 3D turbulence energy does cascade to small scales, the same is not necessarily true in strongly anisotropic 3D flows. Indeed, the dimensional transition typically occurs in two steps: first, as the flow becomes quasi-2D, the forward energy cascade splits up, with some energy cascading to large scales, while 3D variations persist in the flow, transferring the remainder of the energy to small scales. Provided $L\ll L_{\alpha}$, even a small inverse energy flux can lead to condensation at late times, thus drastically altering the flow. Second, for very strong anisotropy, 3D variations decay and exact 2D flow is obtained (forcing and boundary conditions permitting).
Below, we briefly summarize important known results on both of these steps of the dimensional transition from 3D to 2D turbulence in different systems. The following section is inspired by the work of AlexakisAlexakis (2023) with the permission of the author.
@@ III.1 Thin-layer turbulence
where $\overline{(\cdot)}$ represents the vertical average. Measures of chaos based on Lyapunov exponents provide an alternative approach to studying the transition between 2D and quasi-2D turbulenceClark et al. (2021); Ho, Clark, and Berera (2024), but this will not be discussed here.
@@ III.1.1 Homogeneous quasi-2D turbulence
The first study showing that energy can cascade simultaneously to large and small scales in turbulence confined within thin layers is by Smith et al.,Smith, Chasnov, and Waleffe (1996) and a number of more detailed subsequent works later quantified energy fluxes as a function of the layer height.Celani, Musacchio, and Vincenzi (2010); Musacchio and Boffetta (2017); Benavides and Alexakis (2017); van Kan and Alexakis (2019); Poujol, van Kan, and Alexakis (2020) Below, based on these works, we describe how $D_{\nu}^{\mathcal{E}},D_{\alpha}^{\mathcal{E}}$ $\mathcal{E}_{LS}$ and $\mathcal{E}_{3D}$ change as the layer height $H$ is varied.
@@
Figure 3: Fraction of injected energy cascading to large scales for turbulence within thin a layer of height $H$ and horizontal length $L$ forced at scale $\ell_{f}=2\pi/k_{f}$. For sufficiently large scale separation between forcing scale and domain size, the critical scaling of the energy flux is approximately linear. Reproduced with permission from J. Fluid Mech. 822, 364–385 (2017). Copyright 2017 Cambridge University Press.
@@
where an exponent $\beta_{1}$ close to unity is observed, provided a large scale separation between the forcing scale $\ell_{f}$ and the horizontal domain size $L$ (see Fig. 3) but more work is required to measure it more precisely and complement the existing numerical evidence for free slipBenavides and Alexakis (2017) and periodicvan Kan and Alexakis (2019) boundary conditions in the vertical direction. No theory exists to describe the physics of this non-equilibrium transition, and therefore more detailed numerical and theoretical studies will be needed to clarify its properties and determine whether it may potentially be related to a known universality class. Recent results derived using renormalization group methods from field theory indicate the presence of a (fractional) cross-over dimension $d\approx 2.15$ where the energy flux changes sign, but these results are derived based on assumptions such as a spatially infinite domain, by contrast with the finite-size setup discussed here, and more work remains to be done to examine possible connections.Verma (2024) The very existence of a critical point at $H=H_{3D}$ is nontrivial, and remains to be further investigated, given that even bifurcations in extremely well-controlled low-dimensional dynamical systems show some rounding due to experimental imperfections. While the available evidencePoujol, van Kan, and Alexakis (2020) indicates that the existence of $H_{3D}$ is robust, whether the forcing acts on 2D modes alone or a mixture of 2D and 3D modes, more work is needed to further clarify the impact of the forcing dimensionality on this transition.
@@
where $\eta=\epsilon k_{f}^{2}$ is the enstrophy flux, although this remains to be tested in full direct numerical simulations. The energy reaching the scale $H$ is then transported, by 3D interactions, to even smaller scales, where it is dissipated by viscosity.
@@
Figure 4: Visualization of $\omega_{z}^{2}$ in typical flow states of thin-layer turbulence, showing a top view and a side view of the layer. Left panel: Flow in a deep layer at $H>H_{\rm 3D}$ (see main text). Right panel: large-scale vortex condensate in a thin layer at $H<H_{\rm 3D}$. Figure adapted from van Kan et al. (2019)van Kan and Alexakis (2019). Source: Adrian van Kan and Alexandros Alexakis, J. Fluid Mech. 864, 490–518, 2019; licensed under a Creative Commons Attribution (CC BY) license.
@@ III.1.2 The quasi-2D condensate regime
Next, we consider the case where the bottom drag is small or absent. In a finite domain, even a small inverse energy flux then leads to the formation of a large-scale condensate. While numerical simulations of large-scale condensates are computationally very demanding since the saturation of the condensate amplitude is a slow process requiring very long integration times, a few numerical studies exist in the literaturevan Kan and Alexakis (2019); Musacchio and Boffetta (2019) which explicitly address condensate formation in thin-layer turbulence. Laboratory experiments, on the other hand, do not suffer from the same restrictions in terms of observation time as simulations, although they are more limited when it comes to varying the layer geometry or the forcing scale. A number of experimental studies of condensates in thin-layer turbulence have been reported,Xia, Shats, and Falkovich (2009); Shats, Byrne, and Xia (2010); Xia et al. (2011); Byrne, Xia, and Shats (2011); Xia and Francois (2017); Vernet et al. (2021) where the flow is typically driven electromagnetically, although these have not yet been able to explore the layer height dependence in the same detail.
As in the above discussion of the homogeneous turbulence case, we begin by considering a deep layer of height $H\gg\ell_{f}$, where a forward energy cascade is present and no condensate forms. As the layer height is reduced to the critical height $H=H_{3D}$ (discussed above), a weak inverse cascade emerges leading to the formation of a condensate with large-scale energy $\mathcal{E}_{LS}$. It was shownvan Kan and Alexakis (2019) (for the case $\alpha=0$) that the transition to the condensate is discontinuous. Specifically, for $H$ slightly larger than $H_{3D}$, no condensate formed when initializing the flow at rest, and the large-scale $\mathcal{E}_{LS}$ was negligible. By contrast, for $H$ just below $H_{3D}$, the large-scale energy $\mathcal{E}_{LS}$ jumped to a finite value, taking up nearly all of the kinetic energy of the flow, indicating the presence of the condensate. Furthermore, the transition was shown to be hysteretic: as $H$ was increased again, the condensate was found to persist at values of $H$ where it did not spontaneously emerge from small-amplitude initial conditions within the integration time. In other words, there is a coexistence of two attractors (illustrated in Fig. 4 in terms of $\omega_{z}^{2}$) in a range of $H$ near $H_{3D}$. When simulating the system within this parameter range for very long times, one observes rare spontaneous transitions between the two attactors, induced by turbulent fluctuations, occurring at random times. This phenomenon will be described in more detail in section IV.
@@
As in the homogeneous quasi-2D turbulence case, when the layer height is reduced to the order of the Kolmogorov scale, another phase of turbulence is encountered where 3D perturbations decay viscously. However, the dynamics of 3D modes at $H$ slightly above $H_{2D}$ are strongly impacted by the presence of the condensate. In contrast with homogeneous quasi-2D turbulence, where 3D perturbations can develop anywhere in the flow, the condensate flow is rendered highly inhomogeneous by the presence of large-scale coherent structures, with larger strain and vorticity in some regions than others, while the flow still remains chaotic. The evolution of infinitesimal 3D perturbations on the background of a turbulent 2D condensate was recently investigated by Seshasayanan et al. in direct numerical simulations,Seshasayanan and Gallet (2020); Lohani, Nayak, and Seshasayanan (2024) revealing that $\mathcal{E}_{3D}$ in this limit exhibits a random behavior with long periods of slow viscous decay and short periods of rapid exponential increase. The periods of increase were found to appear when extremes of the vorticity or the strain rate crossed a certain threshold. Considering 2D point vortex flow coupled to localized 3D perturbations, van Kan et al.van Kan, Alexakis, and Brachet (2021a) subsequently showed that the instantaneous growth rate of the energy of the 3D perturbations displayed power law distributions that were linked to the power law distribution of strain in space. This indicates that the logarithm of the energy of the 3D perturbations undergoes Lévy random motion, explaining the sudden jumps in the growth of the perturbation energy observed in the simulations of Seshasayanan et al. This led to the investigation of this type of random process in simplified stochastic dynamical systems, where a new type of intermittency was identified (distinct from the dissipative intermittency of 3D turbulence mentioned in Sec. II), so-called Lévy on-off intermittency,van Kan, Alexakis, and Brachet (2021b); Van Kan and Pétrélis (2023) namely, stochastic dynamics near a phase transition subject to multiplicative power-law noise. In particular, this theory predicts a distinct power-law behavior of the 3D energy
@@ III.2 Rapidly rotating turbulence
which was also verified in the limit $\rm Ro\to 0$ using a reduced, asymptotically valid set of equations.van Kan and Alexakis (2020) Figure 5 shows the fraction of kinetic energy cascading upscale as a function of $(k_{f}H\,Ro)^{-1}$ transitions from zero to nonzero values at a critical threshold $(k_{f}H\,{\rm Ro})^{-1}\approx 0.03$, with approximately linear scaling near onset (different symbols indicate different Reynolds numbers and forcing scales).
Figure 5: Fraction of energy cascading upscale in steady state, $D_{\alpha}^{\mathcal{E}}$ as a function of the parameter $(k_{f}H\,Ro)^{-1}$, showing a transition from zero to nonzero values at a threshold value around $(k_{f}H\,Ro)^{-1}\approx 0.03$, with an approximately linear scaling near onset. Different symbols indicate different forcing scales and hyperviscous Reynolds numbers. Reproduced with permission from J. Fluid Mech., 899, A33 (2020). Copyright 2020 Cambridge University Press.
@@
Stably stratified turbulence, or stratified turbulence for short, refers to turbulence under the influence of gravity $-g\hat{\mathbf{z}}$ and an imposed, stable background density gradient $S=-\rho_{0}^{-1}{\rm d}\rho(z)/{\rm d}z>0$, where $\rho_{0}$ is a reference density. Like rotating flow, stratified flow supports wave motion, namely internal waves whose frequency depends on stratification. Unlike rotation, however, stratification acts to suppress the formation of 2D turbulence: by inhibiting vertical motion due to the energy cost associated with lifting up heavy fluid parcels, it leads to the formation of thin layers, creating strong vertical gradients in the process. The strength of stratification is measured by the nondimensional Froude number ${\rm Fr}=\epsilon^{1/3}/(gS)^{1/2}/\ell_{f}^{2/3}$, where small values of $\rm Fr$ indicate strong stratification. Sozza et al.Sozza et al. (2015) showed that the critical height, at which the bidirectional energy cascade emerges, decreases with increasing stratification strength as
@@ III.5 Convection
Convection refers to fluid motion induced by buoyancy in the presence of an imposed unstable background density gradient ($S<0$). The paradigmatic model of this process is Rayleigh-Bénard convection, where the fluid is contained between a hot bottom boundary and a cold top boundary, both of which are typically assumed to be either of no-slip or stress-free type. Convection induces three-dimensional vertical motion. The resulting dynamics are highly nontrivial, in particular since the buoyant forcing is self-organized in response to the flow, rather than being an externally imposed body force. Despite these differences, it has been found that convection, under highly anisotropic conditions, can also display quasi-2D behavior and the formation of large-scale flow structures.
The most well-studied example is that of rapidly rotating convection.Julien et al. (2012a); Ecke and Shishkina (2023) There, for sufficiently small Rossby numbers, an inverse energy cascade forms and results in the formation of large-scale vortices or jets,Guervilly, Hughes, and Jones (2014); Favier, Silvers, and Proctor (2014); Rubio et al. (2014); Guervilly and Hughes (2017); Julien, Knobloch, and Plumley (2018); Maffei et al. (2021) not unlike the systems discussed in the previous sections. Furthermore, the behavior of the large-scale vortex condensates also resembles thin-layer and rotating turbulence, since they emerge via a discontinuous transition.Favier, Guervilly, and Knobloch (2019); de Wit et al. (2022) However, the case of convection has additional complexity, in particular due to the presence of thermal and momentum boundary layers. It is known that the emergence of the inverse cascade is impacted by the boundary conditions. In the presence of stress-free boundaries, at moderately rapid rotation, i.e, moderately small Rossby number and Ekman number $Ek=\sqrt{\nu/(2\Omega H^{2})}$, large-scale vortices form, which are suppressed with no-slip boundaries by viscous Ekman boundary layersKunnen et al. (2016), where Ekman pumping plays an important role.Julien et al. (2016) Nonetheless, at sufficiently rapid rotation rates (sufficiently low $Ek$) the inverse energy cascade and associated large-scale vortices are also observed in the case of no-slip boundaries.Song, Shishkina, and Zhu (2024) Importantly, the corresponding problem of the robustness of inverse cascades and condensation with respect to no-slip boundaries, which are typically present in laboratory experiments, remains to be studied in mechanically forced thin-layer turbulence. Recent progress in numerically simulating convection at very low Rossby numbers and Ekman numbers as low as $Ek=10^{-15}$ has furthermore revealed a phase transition between fully rotationally constrained geostrophic turbulence and a regime where (thermal) boundary layers exhibit ageostrophic turbulence,van Kan et al. (2024a); Julien et al. (2024) which had been predicted theoretically.Julien et al. (2012b)
@@ IV.1 Reversals of mean fields
Spontaneously reversing large-scale fields are commonly found in turbulent fluid flows. An important example is the Earth’s magnetic field, which is thought to arise from the dynamo effect Fauve and Petrelis (2003) associated with motions in the liquid metal outer core, and whose polarity is known to reverse at random times.Jacobs (1994) By contrast, the Sun’s magnetic field follows the more regular 22-year solar cycle.Charbonneau (2014) Magnetohydrodynamic dynamo reversals have also been reproduced in laboratory experiments,Berhanu et al. (2007) and were analyzed extensively at the theoretical level.Pétrélis et al. (2009); Gallet et al. (2012) An additional example of a reversing flow is found at high altitudes in the Earth’s atmosphere, where the so-called quasi-biennial oscillation consists in reversals of the mean zonal wind approximately every two years,Baldwin et al. (2001) a phenomenon that has also been studied using laboratory analogs.Plumb and McEwan (1978); Semin et al. (2018) In addition, reversals are also common in convectively driven flows and have been studied in detail in that context.Sugiyama et al. (2010); Ni, Huang, and Xia (2015); Wang et al. (2018); Chen et al. (2019); Winchester, Dallas, and Howell (2021); Liu, Julien, and Knobloch (2022); Liu et al. (2024)
Another setting where random reversals are observed is that of confined, large-scale, 2D or quasi-2D turbulent flows, which have been studied both in the laboratorySommeria (1986); Michel et al. (2016); Fauve et al. (2017); Pereira, Gissinger, and Fauve (2019) and in theoretical work.Mishra et al. (2015); Shukla, Fauve, and Brachet (2016); Dallas, Seshasayanan, and Fauve (2020a) Shukla et al.Shukla, Fauve, and Brachet (2016) introduced the approach of investigating the statistics of the large-scale modes based on the truncated Euler equation of ideal fluids, retaining only a finite number $N$ of Fourier modes with wave numbers below a chosen cutoff. Such truncations naturally arise in any standard pseudospectral PDE solver. They considered the flow inside a square domain $[0,\pi]^{2}$ with free-slip boundary conditions, where the stream function can be expanded as $\psi(x,y)=\sum_{n,m}\hat{\psi}_{n,m}\sin(nx)\sin(my)$, with the sum being over a truncated set containing $N$ modes. Shukla et al. compared direct numerical simulations of the 2D Navier-Stokes equation for a viscous fluid with corresponding solutions of the 2D truncated Euler equation in the condensate regime and were able show good qualitative (but not quantitative) agreement between the statistics of the largest-scale mode in the two cases, with the appearance of reversals of the large-scale circulation, which become less and less frequent as friction is reduced, leading to a stronger condensate.Similar findings were also made in a forced 2D shear flow.Dallas, Seshasayanan, and Fauve (2020b)
@@
Figure 7: Left panel: Spontaneous reversals seen in the time series of the large-scale mode $\hat{\psi}_{1,1}$ in a confined 2D flow described by the Euler equation truncated to $N=13$ modes, obtained from numerical simulations. Right panel: probability density function (PDF) $p(\hat{\psi}_{1,1})$ versus large-scale circulation $\hat{\psi}_{1,1}$ with different colored lines representing different values of the control parameter $\delta=\Omega/(k_{2}^{2}\mathcal{E})-1$ ($\delta>0$ increases from bottom to top). The black dashed lines indicate the theoretically predicted form of $p(\hat{\psi}_{1,1})$ at different values of $\delta$, which agrees well with the numerical result. Reproduced with permission from Phil. Trans. R. Soc. A. 380:20210049 (2022). Copyright 2022 Royal Society.